### v1

# ACE-RTL: When Agentic Context Evolution Meets RTL-Specialized LLMs

Abstract.
Recent advances in large language models (LLMs) have sparked growing interest in applying them to hardware design automation, particularly for accurate RTL code generation. Prior efforts follow two largely independent paths: (i) training domain-adapted RTL models to internalize hardware semantics, (ii) developing agentic systems that leverage frontier generic LLMs guided by simulation feedback. However, these two paths exhibit complementary strengths and weaknesses. In this work, we present ACE-RTL that unifies both directions through Agentic Context Evolution (ACE). ACE-RTL integrates an RTL-specialized LLM, trained on a large-scale dataset of 1.7 million RTL samples, with a frontier reasoning LLM through three synergistic components: the generator, reflector, and coordinator. These components iteratively refine RTL code toward functional correctness. We further introduce a parallel scaling strategy that significantly reduces the number of iterations required to reach correct solutions. On the Comprehensive Verilog Design Problems (CVDP) benchmark, ACE-RTL achieves up to a $44.87\%$ pass rate improvement over $14$ competitive baselines while requiring only four iterations on average.
copyright: none11footnotetext: Work done during affiliation with NVIDIA, now at Agentrys.
1. Introduction
Large language models (LLMs) have shown rapid progress in software development and are increasingly explored for hardware design automation (Jimenez et al., 2023; Cursor AI, 2025; Anthropic, 2025b; Liu et al., 2023a; Deng et al., 2025a). As hardware systems continue to grow in scale and complexity, developing register-transfer-level (RTL) designs remains a labor-intensive and error-prone process requiring deep domain expertise. Leveraging LLMs to translate natural-language intent into functionally correct RTL code can significantly accelerate design cycles. Moreover, accurate RTL generation is essential for downstream workflows such as performance optimization and architectural design space exploration (Zhang et al., 2025; Yao et al., 2024; Wang et al., 2025).
In parallel to this motivation, the research community has pursued two main directions, as shown in Figures 1(a) and 1(b), for applying LLMs to RTL code generation. (i) The first focuses on developing domain-specialized LLMs trained on RTL data, where models are fine-tuned to internalize RTL syntax and hardware design semantics. For instance, RTLCoder (Liu et al., 2024c, 2025) develops a specialized model fine-tuned on a dedicated Verilog dataset. CraftRTL (Liu et al., 2024b) further enhances RTL coding capability by training on synthetic data that capture non-textual design representations. More recently, ScaleRTL (Deng et al., 2025b) scales up post-training on a large RTL reasoning corpus, yielding a strong specialized model that achieves state-of-the-art (SoTA) results on mainstream RTL benchmarks. (ii) The second direction centers on agentic systems that leverage frontier generic LLMs together with RTL-specific tools such as simulators and waveform analyzers. Along this line, VerilogCoder (Ho et al., 2025) introduces a multi-agent framework based on GPT4-Turbo (OpenAI, 2024) that decomposes RTL design into fine-grained sub-tasks through a task-and-circuit relation graph and repairs functional errors using a waveform tracing tool. Similarly, MAGE (Zhao et al., 2024) employs multiple Claude3.5-Sonnet (Anthropic, 2024) LLM agents tightly integrated with hardware-domain toolchains to produce functionally correct RTL code.
Despite these advances, the two research directions remain largely orthogonal and exhibit complementary limitations. Training-based approaches often yield RTL-specialized models with limited general capabilities such as long-context reasoning, multi-step planning, and instruction following, reducing their effectiveness on complex design tasks. Conversely, agent-based methods lack RTL-specific knowledge learned from large-scale hardware datasets, causing the system to struggle on problems that require deep hardware semantics and domain expertise. These limitations highlight the need for a unified framework that integrates an RTL-specialized model with a frontier general-purpose LLM, enabling the system to leverage the strengths of both paradigms.
Refer to caption Placeholder.
Figure 1. Comparison of prior works and ACE-RTL.
In this work, we advance LLM-based RTL coding by synergizing the complementary strengths of an RTL-specialized model and a frontier general-purpose LLM. To this end, we propose ACE-RTL, which adopts the notion of Agentic Context Evolution (ACE) through three components: the Generator, the Reflector, and the Coordinator. As illustrated in Figure 2(a), the Generator is an RTL-specialized LLM trained on 1.7 million diverse RTL samples. When the Generator produces buggy code, the Reflector, powered by the frontier LLM Claude4-Sonnet (Anthropic, 2025a), is invoked to analyze simulation feedback, trace errors back to faulty logic, and provide high-level fix guidance. The Coordinator then incorporates this guidance into an incrementally evolving context that guides the Generator in the next iteration of code generation. By iteratively cycling among the Generator, Reflector, and Coordinator, ACE-RTL effectively combines domain-specific RTL expertise with strong general-purpose reasoning to produce functionally correct RTL designs. We further introduce a parallel scaling strategy that launches multiple ACE-RTL processes simultaneously and terminates the remaining ones once any process succeeds. This significantly reduces the total number of iterations needed to obtain correct RTL implementations.
To evaluate the efficacy of ACE-RTL, we compare it against the strongest frontier LLMs, as well as SoTA RTL-specialized models. More importantly, instead of relying on prior RTL benchmarks such as VerilogEval (Liu et al., 2023b; Pinckney et al., 2024) and RTLLM (Lu et al., 2024), which primarily consist of relatively simple specification-to-RTL generation tasks (e.g., small modules or textbook-style problems), we conduct evaluations on the Comprehensive Verilog Design Problems (CVDP) benchmark (Pinckney et al., 2025). CVDP encompasses hundreds of diverse and challenging RTL coding tasks, including code completion, specification-to-RTL generation, code modification, and debugging. Since earlier benchmarks have become saturated and no longer sufficiently differentiate recent approaches, we argue that evaluating on CVDP provides a more rigorous and practical assessment of model capability and more convincingly demonstrates meaningful progress in this domain. We summarize our technical contributions as follows:
$\bullet$ To our knowledge, this work is the first to introduce the ACE framework for RTL coding, which jointly leverages a frontier generic LLM and an RTL-specialized model.
$\bullet$ Motivated by data scaling laws, we construct a 1.7M-sample RTL dataset covering diverse coding problems. Trained on this corpus, our RTL-specialized model surpasses SoTA baselines, including GPT-5 and ScaleRTL, by up to $21.45\%$.
$\bullet$ By integrating the RTL-specialized model with the ACE framework, our proposed ACE-RTL achieves up to a $44.87\%$ improvement in pass rate over all baselines on the CVDP benchmark, demonstrating the effectiveness of ACE for complex RTL tasks.
$\bullet$ With parallel scaling, ACE-RTL requires only four iterations on average to produce correct RTL code, achieving up to a $3.06\times$ reduction in total iterations over the non-scaled version.
2. Background
In this section, we review prior studies on RTL-specialized LLMs, agentic systems, and RTL benchmarking, which together establish the foundation and motivation for ACE-RTL.
2.1. Domain-Adapted LLMs for RTL Tasks
Adapting LLMs for RTL coding via model training has attracted growing attention due to their potential to improve hardware design productivity. RTLCoder (Liu et al., 2024c, 2025) develops a GPT-assisted pipeline for automated RTL data synthesis, while CraftRTL (Liu et al., 2024b) generates synthetic, correct-by-construction samples derived from design artifacts. OriGen (Cui et al., 2024b) adopts compiler-guided reflection and augmentation to improve functional correctness. Recently, ScaleRTL (Deng et al., 2025b) scales both training data and test-time compute by fine-tuning on billions of RTL reasoning tokens and applying iterative reasoning extension, achieving SoTA results on VerilogEval (Pinckney et al., 2024) and RTLLM (Lu et al., 2024) benchmarks. These efforts collectively advance the development of RTL-specialized models. However, it remains unclear how to effectively integrate such models into an agentic system that can further push the frontier of LLM-based RTL coding.
2.2. Agentic Systems and Context Evolution
Recent agentic systems such as VerilogCoder (Ho et al., 2025) and MAGE (Zhao et al., 2024) have demonstrated the potential of multi-agent collaboration for RTL code generation. VerilogCoder decomposes RTL design into fine-grained subtasks using a Task and Circuit Relation Graph (TCRG) and employs a waveform-tracing debugger to locate functional errors, while MAGE leverages multiple specialized agents for RTL generation, testbench creation, and debugging through simulator-based feedback. Despite their promising results, both methods depend heavily on complex, tool-specific planning and lack an RTL-specialized model that captures broad hardware coding logic, which makes them difficult to generalize beyond predefined task structures. In this work, we take a different perspective: we argue that the key role of an agentic system is to dynamically construct the right context to guide the RTL generator toward correct solutions. To this end, ACE-RTL iteratively evolves contextual information such as prior attempts and feedback to progressively steer an RTL-specialized model toward functional correctness.
2.3. RTL Benchmarks for LLM Evaluation
Prior mainstream benchmarks such as VerilogEval and RTLLM mostly evaluate small, self-contained RTL problems, emphasizing short specification-to-RTL snippets rather than sustained reasoning over realistic design contexts. RTL-Repo (Allam and Shalan, 2024) takes a step toward practical settings by introducing GitHub-derived projects and prompting models to complete redacted regions, yet it remains limited to code completion and omits broader challenges such as specification-to-RTL generation, code modification, or debugging. We argue that rigorous assessment requires benchmarks stressing both context length and task diversity. Therefore, this work focuses on evaluating RTL methods on CVDP (Pinckney et al., 2025), whose problem statements and target code are several orders of magnitude longer than those in VerilogEval. CVDP includes hundreds of complex RTL problems spanning multiple tasks and hardware domains such as arithmetic units, on-chip protocols, memory hierarchies, DSP kernels, and communication blocks. Empirically, we observe that CVDP better distinguishes model capability and more faithfully reflects the demands of practical RTL design.
3. The ACE-RTL Framework
Refer to caption Placeholder.
Figure 2. Overview of ACE-RTL and its parallel scaling strategy.
Figure 2 provides an overview of ACE-RTL, which integrates an RTL-specialized model and a frontier generic LLM through three synergistic components: the Generator, Reflector, and Coordinator. At a high level, ACE-RTL operates in an iterative agentic loop that progressively refines RTL code toward functional correctness. These three components are discussed in detail in Sections 3.1, 3.2, and 3.3, respectively. In addition, Section 3.4 introduces a parallel scaling strategy that executes multiple ACE-RTL instances concurrently to speedup convergence efficiency while preserving correctness.
3.1. Generator: An RTL-Specialized LLM
We train an RTL-specialized model as ACE-RTL-Generator, which aims to produce correct code, conditioned on an input specification and an evolving context constructed during prior iterations. To this end, we follow two steps to construct a large-scale RTL dataset: (i) collecting and filtering raw RTL scripts from public sources, (ii) generating diverse specification–code pairs via in-context learning.
High-Quality RTL Script Collection
Motivated by data scaling laws in LLMs (Hoffmann et al., 2022; Kaplan et al., 2020), we begin by collecting five million raw RTL scripts from public repositories and open hardware projects. To ensure data quality, we perform a sequence of filtering and validation steps. First, we remove duplicate files and eliminate machine-generated code, including netlist scripts or RTL generated from high-level synthesis. We also filter out extreme outliers by excluding scripts with fewer than $30$ lines or more than $2,000$ lines. Next, we apply syntax validation using the Icarus Verilog (iverilog) (Williams and Baxter, 2002) to discard all code samples containing syntax errors.
To prevent contamination of downstream evaluation benchmarks, we use the Jaccard similarity metric defined in (Deng et al., 2025b) to measure the similarity between each collected RTL snippet and all benchmark golden solutions. Any collected code whose similarity score exceeds $0.8$ with respect to any golden solution is discarded. This effectively prevents overlap with test data and ensures fair evaluation. After all filtering stages, we retain 157K high-quality RTL scripts as the foundation for specification-code pair generation.
Specification and Code Pair Generation
Given the filtered RTL corpus, we employ in-context learning to automatically generate diverse specifications for each RTL design. We first manually construct a pool of 32 high-quality examples, each containing a specification paired with its RTL implementation. These examples span multiple tasks, including RTL code generation, modification, and debugging. For each raw RTL script, we randomly sample one instance from this pool as context and prompt multiple LLMs, including open-source models such as GPT-OSS-120B (Agarwal et al., 2025) and DeepSeek-R1 (Guo et al., 2025) as well as proprietary ones, to produce diverse pairs of task-specific specifications and corresponding golden RTL implementations. Since the generated RTL may deviate from the original design, we apply an additional filtering stage to remove samples with syntax errors or potential benchmark contamination, identified by high similarity to downstream evaluation benchmarks. For modification and debugging tasks, the specification includes a simplified or intentionally faulty version of the golden code, emulating realistic design correction workflows. Through this pipeline, we curate a large-scale dataset of 1.7 million specification–RTL pairs that comprehensively cover diverse hardware design scenarios.
Using this large-scale RTL dataset, we apply supervised fine-tuning (SFT) (Chung et al., 2024) to obtain the RTL-specialized model. The resulting model internalizes hardware domain knowledge and RTL semantics, enabling it to generate correct RTL code across diverse design tasks.
3.2. Reflector: A Frontier Reasoning LLM
While the Generator produces RTL code grounded in domain knowledge, the Reflector analyzes simulation feedback and guides refinement. Frontier generic LLMs excel at interpreting specifications, buggy code, and error reports, making them well suited for this role. In ACE-RTL, the Reflector is built upon Claude4-Sonnet to interpret verifiable feedback and generate structured fix strategies.
At each iteration, the RTL code produced by the Generator is compiled and simulated using iverilog to verify functional correctness. When simulation fails, the Reflector is triggered to analyze the feedback. Automated scripts invoke the simulator and convert the raw logs into a structured format that captures the error message and the expected versus actual signal behavior. Claude4-Sonnet then serves as the Reflector’s reasoning engine: it integrates this parsed information with the design specification and the buggy RTL code to produce a structured diagnostic report. The report provides a concise explanation of the root cause together with high-level fix guidance describing how to resolve the primary failure. This structured output is then passed to the Coordinator for context evolution. Notably, the entire process depends only on iverilog, a lightweight and open-source simulator, rather than complex proprietary toolchains. This design choice makes the system portable, easy to maintain, and readily adaptable to diverse RTL design tasks.
3.3. Coordinator: A Context Evolution Engine
The Coordinator in ACE-RTL serves as the memory and control center of the agentic loop, maintaining a dynamic, self-evolving context that guides the Generator in subsequent iterations. Its primary role is to provide an incrementally refined context that captures an organized record of prior attempts, enabling more informed reasoning while avoiding redundant or counterproductive fixes. This component is implemented through efficient scripts, with Claude4-Sonnet providing core reasoning support, to continuously update and restructure the evolving context based on the system’s progress.
The Coordinator performs two key functions. First, it aggregates debugging history across iterations, tracking which errors were identified, what fixes were suggested, and whether those fixes resolved the issues. This structured memory preserves useful progress and prevents regressions, allowing the Generator to focus on remaining failure cases rather than revisiting already-corrected logic.
Second, the Coordinator monitors progress across iterations. If the same error persists over multiple rounds, it recognizes that the current code trajectory may be difficult to repair and triggers a restart mechanism, discarding the current implementation and prompting the Generator to produce a fresh RTL design from the specification. This leverages LLM stochasticity to explore alternative implementations that is easier to correct.
By combining structured history aggregation with an adaptive restart mechanism to construct a self-evolving context, the Coordinator keeps ACE-RTL both stable and flexible. It preserves useful debugging insights when progress is made and triggers a controlled reset when stagnation occurs, enabling more efficient progress toward functionally correct RTL code.
3.4. Parallel Scaling in ACE-RTL
Although the iterative debugging framework in ACE-RTL effectively improves code correctness, it may require a large number of iterations to converge when starting from an unfavorable initial RTL generation. To address this inefficiency, we introduce a parallel scaling strategy that exploits the inherent stochasticity of LLM-based code generation to accelerate convergence.
Instead of relying on a single debugging trajectory, the parallel scaling strategy launches multiple ACE-RTL processes simultaneously, as shown in Figure 2(b). Each process begins with a distinct RTL implementation generated by the Generator from the same input specification. These processes then proceed separately through the iterative loop of the Generator, Reflector, and Coordinator components, performing debugging and refinement in parallel.
During execution, all processes are monitored concurrently. Once any process produces an RTL implementation that passes all functional test cases, the system immediately terminates all remaining processes and outputs the correct code. If none of the processes achieve functional correctness within the predefined maximum number of iterations, all are stopped automatically.
This strategy leverages the diversity introduced by LLM stochasticity, as some initial generations naturally require far fewer debugging steps than others. By exploring multiple debugging paths in parallel, the system reduces the worst-case iterations needed to reach a correct solution. Consequently, the parallel scaling strategy largely reduces the overall iteration count, as confirmed in Figure 6.
4. Experiment
We conduct a comprehensive evaluation of ACE-RTL, comparing its performance against $14$ competitive baselines on the CVDP benchmark. In addition, we present three case studies illustrating the effectiveness of key components within ACE-RTL. Besides, we also demonstrate how the proposed parallel scaling strategy largely reduces the number of iterations required to reach correct solutions.
4.1. Experimental Setup
Baselines
We evaluate ACE-RTL against $13$ SoTA baselines, including latest large open-source models Llama4-Maverick (AI, 2024), DeepSeek-v3.1 (Liu et al., 2024a), DeepSeek-R1 (Guo et al., 2025), Kimi-K2 (Team et al., 2025), and Qwen3-Coder-480B (Yang et al., 2025). Frontier proprietary models comprise 4o-mini (Hurst et al., 2024), GPT-5 (OpenAI, 2025), and Claude4-Sonnet (Anthropic, 2025a). For RTL models, we include RTLCoder-v1.1, CodeV (Zhao et al., 2025), OriGen (Cui et al., 2024a), CraftRTL, and ScaleRTL. Besides, given that ScaleRTL extends ScaleRTL via test-time scaling and achieves performance comparable to prior RTL agentic methods such as VerilogCoder and MAGE, we select it as the representative agentic baseline for evaluation on CVDP. Our variants are: (1) ACE-RTL-Generator, our RTL-specialized model, and (2) ACE-RTL, which integrates all components of this work. (3) ACE-RTL (Claude4), which uses the same agentic framework as ACE-RTL but replaces our specialized Generator with Claude4-Sonnet.
CVDP Benchmark
We adopt the CVDP-v1.0.2 benchmark as our primary evaluation suite, which includes four representative RTL design tasks: code completion (CVDP-cid002), specification-to-RTL generation (CVDP-cid003), code modification (CVDP-cid004), and code debugging (CVDP-cid016). For the latter two tasks, the input prompt includes a basic or buggy RTL implementation that is intended to be refined according to the given specification. Together, these tasks capture the core challenges faced by RTL designers in practice, covering both generative and corrective design scenarios. The benchmark contains $94$, $78$, $55$, and $35$ problems under the four categories respectively, covering a diverse range of hardware domains including arithmetic units, on-chip communication protocols, memory hierarchies, DSP kernels, and control logic. We leave testbench generation and subjective evaluation categories in CVDP to our future work. Each problem is evaluated across five independent runs. We report two complementary metrics: (1) Pass@1, the average success rate across independent runs (Liu et al., 2023b), and (2) Agentic Pass Rate (APR), defined as the ratio of unique solved problems to the total number of problems. Since agentic methods inherently perform multiple reasoning and correction iterations, comparing them against standalone LLMs solely using Pass@1 can be misleading. Therefore, we introduce APR to better capture a model’s overall problem-solving coverage across multiple trials, providing a fairer assessment of agentic approaches.
Training and Inference
ACE-RTL-Generator is finetuned from Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) using 32 compute nodes, each equipped with 8 NVIDIA A100 GPUs. We adopt tensor, pipeline, and context parallelism to enable efficient large-scale fine-tuning. Training is performed for three epochs with a context window of 32,76832{,}768 tokens and a global batch size of $128$, using a cosine-annealing learning-rate scheduler. For inference, ACE-RTL uses five parallel processes under our parallel scaling strategy, with each process limited to 30 iterations and the Generator configured with a temperature of $1.2$. The finetuned ACE-RTL-Generator is hosted via vLLM for low-latency API inference (Kwon et al., 2023), while other components leverage Claude4-Sonnet through its official API calls.
Table 1. Functional correctness on the CVDP benchmark. Each benchmark reports percentage scores for Pass@1 and Agentic Pass Rate (APR). APR is defined as the ratio of unique solved problems to the total number of problems. For agentic methods, we omit Pass@1 scores (shown as "—") because their Pass@1 equals their APR. We use bold to denote the best results achieved by our methods and underline to mark the strongest baseline results within each CVDP task category.
Type Model CVDP-cid002 Code Completion CVDP-cid003 Spec-to-RTL CVDP-cid004 Code Modification CVDP-cid016 Code Debugging
Pass@1 APR Pass@1 APR Pass@1 APR Pass@1 APR
Open-source models Llama4-Maverick $26.81$ $28.72$ $29.49$ $32.05$ $36.36$ $38.18$ $36.00$ $37.14$
DeepSeek-v3.1 $32.34$ $37.23$ $41.79$ $48.72$ $36.73$ $41.82$ $34.86$ $40.00$
DeepSeek-R1 $34.89$ $\underline{39.36}$ $39.23$ $42.31$ $37.45$ $43.64$ $48.57$ $51.43$
Kimi-K2 $23.40$ $25.53$ $26.67$ $29.49$ $29.09$ $32.73$ $29.71$ $31.43$
Qwen3-Coder-480B $30.43$ $31.91$ $33.33$ $35.90$ $35.27$ $41.82$ $39.43$ $42.86$
Proprietary models o4-mini $35.10$ $37.23$ $41.56$ $45.45$ $41.48$ $44.44$ $50.00$ $58.82$
GPT-5 $36.17$ $\underline{39.36}$ $42.31$ $47.44$ $\underline{43.64}$ $45.45$ $\underline{54.28}$ $\underline{60.00}$
Claude4-Sonnet $\underline{37.94}$ $\underline{39.36}$ $\underline{49.49}$ $\underline{51.28}$ $42.91$ $\underline{49.09}$ $51.43$ $54.29$
RTL-specialized models RTLCoder-v1.1-7B $0.43$ $1.06$ $3.33$ $5.13$ $1.09$ $1.82$ $0.57$ $2.86$
CodeV-7B $3.83$ $6.38$ $5.38$ $7.69$ $0.00$ $0.00$ $0.00$ $0.00$
OriGen-7B $18.30$ $21.28$ $18.97$ $21.79$ $14.18$ $16.36$ $6.86$ $11.43$
CraftRTL-15B $8.09$ $11.70$ $12.31$ $17.95$ $13.45$ $16.36$ $5.14$ $8.57$
ScaleRTL-32B $25.32$ $27.66$ $28.97$ $33.33$ $25.82$ $30.91$ $30.86$ $37.14$
ScaleRTL-32B $29.79$ $35.90$ $32.73$ $40.00$
Ours ACE-RTL-Generator $\mathbf{39.57}$ $40.43$ $\mathbf{49.74}$ $52.56$ $\mathbf{65.09}$ $67.27$ $\mathbf{56.00}$ $57.14$
ACE-RTL (Claude4) $\mathbf{80.85}$ $89.74$ $81.82$ $88.57$
ACE-RTL $\mathbf{80.85}$ $\mathbf{96.15}$ $\mathbf{90.91}$ $\mathbf{91.43}$
4.2. Main Results
As shown in Table 1, most prior RTL-specialized models struggle on CVDP, despite their strong performance on earlier RTL benchmarks. This suggests that models trained mainly on shorter and simpler RTL data generalize poorly to the more complex and realistic design challenges posed by CVDP. In contrast, recent frontier proprietary models such as GPT-5 and Claude4-Sonnet show notable improvements and set strong new baselines, benefiting from broader RTL exposure during large-scale pretraining and stronger reasoning capabilities on complex coding tasks.
Our RTL-specialized model, ACE-RTL-Generator, further advances the RTL coding frontier, outperforming all baselines on most tasks. It achieves a Pass@1 of $65.09\%$ on Code Modification (cid004), surpassing the SoTA baseline GPT-5 by $21.45\%$. This strong performance stems from our high-quality modification data, created by extracting basic RTL code segments and training the model to augment them with missing features, as discussed in Section 3.1.
By integrating our specialized Generator with the full ACE framework, ACE-RTL achieves the best overall performance on CVDP, with up to a $44.87\%$ APR improvement over SoTA baselines. ACE-RTL also outperforms the ACE-RTL (Claude4) variant, showing that an RTL-specialized model is more effective than relying solely on a frontier generic LLM. Together, these results validate the effectiveness of ACE-RTL, which iteratively evolves contextual information to guide the specialized model toward accurate RTL coding.
4.3. Case Studies
To further demonstrate the effectiveness of our approach, we present three representative case studies. The first case showcases how our RTL-specialized model (ACE-RTL-Generator) exhibits a deeper grasp of hardware design principles and canonical RTL patterns compared to generic LLMs. The second case illustrates how ACE-RTL-Reflector elevates repeated local failures into functional reasoning, enabling it to diagnose structural errors. The third case demonstrates how ACE-RTL-Coordinator identifies unproductive progress and strategically triggers restarts, enabling the system to escape local minima and reach correct RTL implementations. Notably, the examples in the latter two case studies require the highest number of iterations to reach correct solutions in our experiments.
Refer to caption Placeholder.
Figure 3. Case Study I — RS232 Transmitter Module.
Case Study I — This case study examines the RS232 Transmitter module, which performs serial data transmission driven by precise baud-rate timing. A central component is the baud-rate generator, which produces periodic pulses that step the transmitter state machine through start, data, and stop bits. Correct behavior relies on an accurate fractional clock divider to handle the non-integer ratio between the system clock and the target baud rate.
As shown in Figure 3, Claude4-Sonnet fails to produce a functionally correct implementation due to misunderstanding a fundamental RTL design pattern. Its code couples the accumulator and pulse into a single assignment, thereby losing the overflow-based pulse signal and producing unstable timing behavior. In contrast, our RTL-specialized model (i.e., ACE-RTL-Generator) correctly implements an accumulator-based divider, updating only the lower bits and using the most significant bit as a stable pulse output. This solution reflects an understanding of canonical hardware timing structures, which our model learns from curated RTL training data.
Placeholder.
Figure 4. Case Study II.
Placeholder.
Figure 5. Case Study III.
Placeholder.
Figure 6. Effect of parallel scaling.
     Case Study II — This case study analyzes the 64b/66b Decoder module, which performs bit-aligned extraction of mixed-mode data streams. Figure 6 shows that ACE-RTL quickly improves functional correctness but remains stuck for many iterations before passing the final test case. ACE-RTL-Reflector identifies this stagnation and analyzes the discrepancies between expected and actual outputs. By comparing these discrepancies against the provided specification, the Reflector infers that the decoding logic requires an implicit alignment transformation that is not stated in the specification. It then formulates high-level guidance describing this missing behavioral requirement, which directs the Generator to update the extraction logic accordingly, after which the design immediately passes all remaining tests.
This case illustrates how ACE-RTL-Reflector interprets systematic failure signals to detect specification gaps and elevate repeated local adjustments into understanding the implied design intent, enabling the system to resolve design-level issues that cannot be addressed through isolated fixes.
Case Study III — This case study analyzes the Clock Jitter Detection module, which evaluates clock stability by comparing measured edge intervals against a timing threshold. As shown in Figure 6, ACE-RTL makes early progress but soon enters a long plateau where all attempts fail. The stagnation occurs because each iteration repeats the same flawed behavior: performing jitter detection before a valid baseline interval is established. Although the Generator produces many syntactic variants, they all retain this early-trigger issue and therefore cannot satisfy the required timing semantics.
ACE-RTL-Coordinator recognizes this stalled trajectory by observing that many consecutive iterations exhibit the same assertion pattern with no meaningful improvement. Consequently, the Coordinator triggers a restart and distills high-level insights from the failed attempts, such as the need to validate the first measurement interval before performing detection. It then instructs the Generator to regenerate the design from the specification using these clarified constraints.
The restart produces a simpler and more interpretable implementation that shifts the logic toward proper interval validation. After two restarts, the Generator introduces a validity flag that delays jitter detection until the first full measurement interval has been observed, which immediately restores correctness. This case shows how ACE-RTL-Coordinator identifies prolonged stagnation, extracts useful design insights, and uses restart mechanism to move the system toward correct RTL implementations.
4.4. Efficacy of Parallel Scaling
Figure 6 illustrates the impact of the proposed parallel scaling strategy on reducing the iteration count required for convergence across multiple CVDP categories. The results clearly show that launching multiple independent ACE-RTL instances in parallel significantly accelerates the debugging process without compromising correctness. Specifically, the average iteration count drops from $11.33$ to $3.95$ for cid002, from $11.25$ to $4.23$ for cid003, from $9.25$ to $3.75$ for cid004, and from $13.36$ to $4.37$ for cid016, yielding speedups of $2.87\times$, $2.66\times$, $2.47\times$, and $3.06\times$, respectively.
This improvement demonstrates the efficacy of the parallel scaling design: by exploring diverse debugging trajectories derived from different initial generations, ACE-RTL avoids being constrained by any single unfavorable starting point. Instead, at least one parallel process typically converges rapidly, allowing other trajectories to be terminated early. Consequently, the system achieves nearly a $3\times$ runtime reduction to explore a correct solution.
4.5. Results on VerilogEval
Table 2. Pass rate results on VerilogEval-Human-v2 — ACE-RTL-G denotes the standalone ACE-RTL-Generator model.
Pass@1 APR
CraftRTL Claude4-Sonnet ACE-RTL-G ScaleRTL VerilogCoder ACE-RTL
$68.0$ $73.0$ $\mathbf{73.8}$ $93.6$ $94.2$ $\mathbf{95.5}$
For completeness, we also evaluate our approach on VerilogEval-Human-v2. To ensure fair and rigorous comparison, we use Pass@1 for standalone models and APR for agentic methods. As shown in Table 2, ACE-RTL-G outperforms both the strong RTL model CraftRTL and the frontier model Claude4-Sonnet, demonstrating the strength of our RTL-specialized generator. Moreover, ACE-RTL achieves the highest pass rate among all agentic baselines, surpassing ScaleRTL and VerilogCoder. These results show that ACE-RTL is not overfitted to CVDP but instead captures transferable RTL reasoning and coding capabilities, maintaining strong performance on earlier RTL benchmarks.
5. Conclusion
This work introduces ACE-RTL, a framework that integrates an RTL-specialized model with ACE for accurate RTL code generation. ACE-RTL achieves SoTA results on multiple RTL benchmarks, surpassing both frontier proprietary and RTL-specialized models.

### v2

@@
 ACE-RTL: When Agentic Context Evolution Meets RTL-Specialized LLMs
-Abstract.
-Recent advances in large language models (LLMs) have sparked growing interest in applying them to hardware design automation, particularly for accurate RTL code generation.
+Abstract
+Recent advances in LLMs have sparked growing interest in applying them to hardware design automation, particularly for accurate RTL code generation.
 Prior efforts follow two largely independent paths: (i) training domain-adapted RTL models to internalize hardware semantics, (ii) developing agentic systems that leverage frontier generic LLMs guided by simulation feedback.
@@
 These components iteratively refine RTL code toward functional correctness.
-We further introduce a parallel scaling strategy that significantly reduces the number of iterations required to reach correct solutions.
-On the Comprehensive Verilog Design Problems (CVDP) benchmark, ACE-RTL achieves up to a $44.87\%$ pass rate improvement over $14$ competitive baselines while requiring only four iterations on average.
-copyright: none11footnotetext: Work done during affiliation with NVIDIA, now at Agentrys.
-1. Introduction
-Large language models (LLMs) have shown rapid progress in software development and are increasingly explored for hardware design automation (Jimenez et al., 2023; Cursor AI, 2025; Anthropic, 2025b; Liu et al., 2023a; Deng et al., 2025a).
+We further analyze a parallel scaling strategy that reduces wall-clock iterations to first success by exploring diverse debugging trajectories concurrently.
+On the CVDP benchmark, ACE-RTL achieves up to a $41.02\%$ pass rate improvement over $14$ competitive baselines.
+11footnotetext: Work done during affiliation with NVIDIA, now at Agentrys.
+I Introduction
+Large language models (LLMs) have shown rapid progress in software development and are increasingly explored for hardware design automation [16, 8, 5, 19, 9].
 As hardware systems continue to grow in scale and complexity, developing register-transfer-level (RTL) designs remains a labor-intensive and error-prone process requiring deep domain expertise.
 Leveraging LLMs to translate natural-language intent into functionally correct RTL code can significantly accelerate design cycles.
-Moreover, accurate RTL generation is essential for downstream workflows such as performance optimization and architectural design space exploration (Zhang et al., 2025; Yao et al., 2024; Wang et al., 2025).
+Moreover, accurate RTL generation is essential for downstream workflows such as performance optimization and architectural design space exploration [35, 34, 30].
 In parallel to this motivation, the research community has pursued two main directions, as shown in Figures 1(a) and 1(b), for applying LLMs to RTL code generation.
 (i) The first focuses on developing domain-specialized LLMs trained on RTL data, where models are fine-tuned to internalize RTL syntax and hardware design semantics.
-For instance, RTLCoder (Liu et al., 2024c, 2025) develops a specialized model fine-tuned on a dedicated Verilog dataset.
-CraftRTL (Liu et al., 2024b) further enhances RTL coding capability by training on synthetic data that capture non-textual design representations.
-More recently, ScaleRTL (Deng et al., 2025b) scales up post-training on a large RTL reasoning corpus, yielding a strong specialized model that achieves state-of-the-art (SoTA) results on mainstream RTL benchmarks.
+For instance, RTLCoder [23, 22] develops a specialized model fine-tuned on a dedicated Verilog dataset.
+CraftRTL [21] further enhances RTL coding capability by training on synthetic data that capture non-textual design representations.
+More recently, ScaleRTL [10] scales up post-training on a large RTL reasoning corpus, yielding a strong specialized model that achieves state-of-the-art (SoTA) results on mainstream RTL benchmarks.
 (ii) The second direction centers on agentic systems that leverage frontier generic LLMs together with RTL-specific tools such as simulators and waveform analyzers.
-Along this line, VerilogCoder (Ho et al., 2025) introduces a multi-agent framework based on GPT4-Turbo (OpenAI, 2024) that decomposes RTL design into fine-grained sub-tasks through a task-and-circuit relation graph and repairs functional errors using a waveform tracing tool.
-Similarly, MAGE (Zhao et al., 2024) employs multiple Claude3.5-Sonnet (Anthropic, 2024) LLM agents tightly integrated with hardware-domain toolchains to produce functionally correct RTL code.
+Along this line, VerilogCoder [13] introduces a multi-agent framework based on GPT4-Turbo [25] that decomposes RTL design into fine-grained sub-tasks through a task-and-circuit relation graph and repairs functional errors using a waveform tracing tool.
+Similarly, MAGE [37] employs multiple Claude [3] LLM agents tightly integrated with hardware-domain toolchains to produce functionally correct RTL code.
 Despite these advances, the two research directions remain largely orthogonal and exhibit complementary limitations.
-Training-based approaches often yield RTL-specialized models with limited general capabilities such as long-context reasoning, multi-step planning, and instruction following, reducing their effectiveness on complex design tasks.
-Conversely, agent-based methods lack RTL-specific knowledge learned from large-scale hardware datasets, causing the system to struggle on problems that require deep hardware semantics and domain expertise.
+Training-based approaches often yield RTL-specialized models with limited general capabilities such as long-context reasoning, multi-step planning, and instruction following, reducing their efficacy on complex design tasks.
+Conversely, agent-based methods lack hardware knowledge learned from large-scale RTL datasets, causing the system to struggle on problems that require deep hardware semantics and domain expertise.
 These limitations highlight the need for a unified framework that integrates an RTL-specialized model with a frontier general-purpose LLM, enabling the system to leverage the strengths of both paradigms.
-Refer to caption Placeholder.
-Figure 1. Comparison of prior works and ACE-RTL.
+Figure 1: Comparison of prior works and ACE-RTL.
 In this work, we advance LLM-based RTL coding by synergizing the complementary strengths of an RTL-specialized model and a frontier general-purpose LLM.
@@
 As illustrated in Figure 2(a), the Generator is an RTL-specialized LLM trained on 1.7 million diverse RTL samples.
-When the Generator produces buggy code, the Reflector, powered by the frontier LLM Claude4-Sonnet (Anthropic, 2025a), is invoked to analyze simulation feedback, trace errors back to faulty logic, and provide high-level fix guidance.
+Different from prior RTL models that are trained exclusively on specification-to-RTL data, our dataset additionally includes code editing and debugging tasks.
+This design better prepares the Generator for its role in the agentic loop, where it not only generates initial RTL code from specification, but also edits and fixes bugs in code in later iterations.
+When errors occur, the Reflector, powered by the frontier LLM Claude4-Sonnet [4], analyzes simulation feedback, traces errors to faulty logic, and provides high-level fix guidance.
 The Coordinator then incorporates this guidance into an incrementally evolving context that guides the Generator in the next iteration of code generation.
@@
 We further introduce a parallel scaling strategy that launches multiple ACE-RTL processes simultaneously and terminates the remaining ones once any process succeeds.
-This significantly reduces the total number of iterations needed to obtain correct RTL implementations.
-To evaluate the efficacy of ACE-RTL, we compare it against the strongest frontier LLMs, as well as SoTA RTL-specialized models.
-More importantly, instead of relying on prior RTL benchmarks such as VerilogEval (Liu et al., 2023b; Pinckney et al., 2024) and RTLLM (Lu et al., 2024), which primarily consist of relatively simple specification-to-RTL generation tasks (e.g., small modules or textbook-style problems), we conduct evaluations on the Comprehensive Verilog Design Problems (CVDP) benchmark (Pinckney et al., 2025).
-CVDP encompasses hundreds of diverse and challenging RTL coding tasks, including code completion, specification-to-RTL generation, code modification, and debugging.
+This largely reduces the iterations required to reach a correct RTL implementation, with only a modest increase in token usage.
+To evaluate the efficacy of ACE-RTL, we compare it against the strongest frontier LLMs, as well as SoTA RTL-specialized models and agents.
+More importantly, instead of relying on prior RTL benchmarks such as VerilogEval [20, 27] and RTLLM [24], which primarily consist of relatively simple specification-to-RTL generation tasks (e.g., small modules or textbook-style problems), we conduct evaluations on the Comprehensive Verilog Design Problems (CVDP) benchmark [28].
+CVDP encompasses hundreds of diverse and challenging RTL coding tasks.
 Since earlier benchmarks have become saturated and no longer sufficiently differentiate recent approaches, we argue that evaluating on CVDP provides a more rigorous and practical assessment of model capability and more convincingly demonstrates meaningful progress in this domain.
 We summarize our technical contributions as follows:
-$\bullet$ To our knowledge, this work is the first to introduce the ACE framework for RTL coding, which jointly leverages a frontier generic LLM and an RTL-specialized model.
-$\bullet$ Motivated by data scaling laws, we construct a 1.7M-sample RTL dataset covering diverse coding problems.
-Trained on this corpus, our RTL-specialized model surpasses SoTA baselines, including GPT-5 and ScaleRTL, by up to $21.45\%$.
-$\bullet$ By integrating the RTL-specialized model with the ACE framework, our proposed ACE-RTL achieves up to a $44.87\%$ improvement in pass rate over all baselines on the CVDP benchmark, demonstrating the effectiveness of ACE for complex RTL tasks.
-$\bullet$ With parallel scaling, ACE-RTL requires only four iterations on average to produce correct RTL code, achieving up to a $3.06\times$ reduction in total iterations over the non-scaled version.
-2. Background
-In this section, we review prior studies on RTL-specialized LLMs, agentic systems, and RTL benchmarking, which together establish the foundation and motivation for ACE-RTL.
-2.1. Domain-Adapted LLMs for RTL Tasks
+$\bullet$ To our knowledge, ACE-RTL is the first approach to integrate an RTL-specialized model into an agentic system through iterative context evolution, demonstrating that a finetuned RTL model can contribute not only as a strong standalone generator, but also as an effective core component in iterative agentic refinement.
+$\bullet$ Trained on a 1.7M-sample RTL corpus spanning diverse tasks, our RTL model achieves up to a $21.45\%$ higher pass rate than strong standalone baselines such as Claude4 and GPT5, while also aligning well with the iterative bug-fixing setting used in our agentic loop.
+$\bullet$ By combining this RTL model with the ACE framework, ACE-RTL achieves up to a $41.02\%$ pass-rate improvement over all baselines on the CVDP benchmark, demonstrating the effectiveness of iterative context evolution for complex RTL tasks.
+$\bullet$ With parallel scaling, ACE-RTL speeds up convergence to correct RTL code, achieving a $2.77\times$ reduction in wall-clock iterations relative to the non-scaled version, with only a modest increase in token usage.
+II Background
+II-A Domain-Adapted LLMs for RTL Tasks
 Adapting LLMs for RTL coding via model training has attracted growing attention due to their potential to improve hardware design productivity.
-RTLCoder (Liu et al., 2024c, 2025) develops a GPT-assisted pipeline for automated RTL data synthesis, while CraftRTL (Liu et al., 2024b) generates synthetic, correct-by-construction samples derived from design artifacts.
-OriGen (Cui et al., 2024b) adopts compiler-guided reflection and augmentation to improve functional correctness.
-Recently, ScaleRTL (Deng et al., 2025b) scales both training data and test-time compute by fine-tuning on billions of RTL reasoning tokens and applying iterative reasoning extension, achieving SoTA results on VerilogEval (Pinckney et al., 2024) and RTLLM (Lu et al., 2024) benchmarks.
+RTLCoder [23, 22] develops a GPT-assisted pipeline for automated RTL data synthesis, while CraftRTL [21] generates synthetic, correct-by-construction samples derived from design artifacts.
+OriGen [7] adopts compiler-guided reflection and augmentation to improve functional correctness.
+Recently, ScaleRTL [10] scales both training data and test-time compute by fine-tuning on billions of RTL reasoning tokens and applying iterative reasoning extension, achieving SoTA results on VerilogEval [27] and RTLLM [24] benchmarks.
 These efforts collectively advance the development of RTL-specialized models.
-However, it remains unclear how to effectively integrate such models into an agentic system that can further push the frontier of LLM-based RTL coding.
-2.2. Agentic Systems and Context Evolution
-Recent agentic systems such as VerilogCoder (Ho et al., 2025) and MAGE (Zhao et al., 2024) have demonstrated the potential of multi-agent collaboration for RTL code generation.
+However, those prior models are trained predominantly on specification-to-RTL generation data, with limited coverage of code editing and debugging behaviors.
+This training mismatch makes them less effective in practical agentic settings, where the model needs to iteratively repair buggy RTL produced in earlier iterations.
+To tackle this issue, we curate a 1.7M-sample RTL corpus covering code generation, editing, and debugging tasks.
+Training on this corpus yields an RTL-specialized model that is not only strong at standalone RTL generation, but also fits agentic workflows for iterative bug fixing.
+II-B Agentic Systems and Context Evolution
+Recent agentic systems such as VerilogCoder [13] and MAGE [37] have demonstrated the potential of multi-agent collaboration for RTL code generation.
 VerilogCoder decomposes RTL design into fine-grained subtasks using a Task and Circuit Relation Graph (TCRG) and employs a waveform-tracing debugger to locate functional errors, while MAGE leverages multiple specialized agents for RTL generation, testbench creation, and debugging through simulator-based feedback.
-Despite their promising results, both methods depend heavily on complex, tool-specific planning and lack an RTL-specialized model that captures broad hardware coding logic, which makes them difficult to generalize beyond predefined task structures.
+Despite their promising results, both methods depend heavily on complex, tool-specific planning and lack an RTL-specialized model that captures broad hardware coding logic, which can make them harder to generalize beyond predefined task structures.
 In this work, we take a different perspective: we argue that the key role of an agentic system is to dynamically construct the right context to guide the RTL generator toward correct solutions.
-To this end, ACE-RTL iteratively evolves contextual information such as prior attempts and feedback to progressively steer an RTL-specialized model toward functional correctness.
-2.3. RTL Benchmarks for LLM Evaluation
+To this end, ACE-RTL iteratively evolves contextual information from previous iterations to progressively steer an RTL model toward functional correctness.
+II-C RTL Benchmarks for LLM Evaluation
 Prior mainstream benchmarks such as VerilogEval and RTLLM mostly evaluate small, self-contained RTL problems, emphasizing short specification-to-RTL snippets rather than sustained reasoning over realistic design contexts.
-RTL-Repo (Allam and Shalan, 2024) takes a step toward practical settings by introducing GitHub-derived projects and prompting models to complete redacted regions, yet it remains limited to code completion and omits broader challenges such as specification-to-RTL generation, code modification, or debugging.
+RTL-Repo [2] takes a step toward practical settings by introducing GitHub-derived projects and prompting models to complete redacted regions, yet it remains limited to code completion and omits broader challenges such as specification-to-RTL generation, code modification, or debugging.
 We argue that rigorous assessment requires benchmarks stressing both context length and task diversity.
-Therefore, this work focuses on evaluating RTL methods on CVDP (Pinckney et al., 2025), whose problem statements and target code are several orders of magnitude longer than those in VerilogEval.
+Therefore, this work focuses on evaluating RTL methods on CVDP [28], whose problem statements and target code are several orders of magnitude longer than those in VerilogEval.
 CVDP includes hundreds of complex RTL problems spanning multiple tasks and hardware domains such as arithmetic units, on-chip protocols, memory hierarchies, DSP kernels, and communication blocks.
 Empirically, we observe that CVDP better distinguishes model capability and more faithfully reflects the demands of practical RTL design.
-3. The ACE-RTL Framework
-Refer to caption Placeholder.
-Figure 2. Overview of ACE-RTL and its parallel scaling strategy.
+III The ACE-RTL Framework
+Figure 2: Overview of ACE-RTL and its parallel scaling strategy.
 Figure 2 provides an overview of ACE-RTL, which integrates an RTL-specialized model and a frontier generic LLM through three synergistic components: the Generator, Reflector, and Coordinator.
 At a high level, ACE-RTL operates in an iterative agentic loop that progressively refines RTL code toward functional correctness.
-These three components are discussed in detail in Sections 3.1, 3.2, and 3.3, respectively.
-In addition, Section 3.4 introduces a parallel scaling strategy that executes multiple ACE-RTL instances concurrently to speedup convergence efficiency while preserving correctness.
-3.1. Generator: An RTL-Specialized LLM
-We train an RTL-specialized model as ACE-RTL-Generator, which aims to produce correct code, conditioned on an input specification and an evolving context constructed during prior iterations.
-To this end, we follow two steps to construct a large-scale RTL dataset: (i) collecting and filtering raw RTL scripts from public sources, (ii) generating diverse specification–code pairs via in-context learning.
-High-Quality RTL Script Collection
-Motivated by data scaling laws in LLMs (Hoffmann et al., 2022; Kaplan et al., 2020), we begin by collecting five million raw RTL scripts from public repositories and open hardware projects.
-To ensure data quality, we perform a sequence of filtering and validation steps.
-First, we remove duplicate files and eliminate machine-generated code, including netlist scripts or RTL generated from high-level synthesis.
-We also filter out extreme outliers by excluding scripts with fewer than $30$ lines or more than $2,000$ lines.
-Next, we apply syntax validation using the Icarus Verilog (iverilog) (Williams and Baxter, 2002) to discard all code samples containing syntax errors.
-To prevent contamination of downstream evaluation benchmarks, we use the Jaccard similarity metric defined in (Deng et al., 2025b) to measure the similarity between each collected RTL snippet and all benchmark golden solutions.
-Any collected code whose similarity score exceeds $0.8$ with respect to any golden solution is discarded.
-This effectively prevents overlap with test data and ensures fair evaluation.
-After all filtering stages, we retain 157K high-quality RTL scripts as the foundation for specification-code pair generation.
-Specification and Code Pair Generation
-Given the filtered RTL corpus, we employ in-context learning to automatically generate diverse specifications for each RTL design.
-We first manually construct a pool of 32 high-quality examples, each containing a specification paired with its RTL implementation.
-These examples span multiple tasks, including RTL code generation, modification, and debugging.
-For each raw RTL script, we randomly sample one instance from this pool as context and prompt multiple LLMs, including open-source models such as GPT-OSS-120B (Agarwal et al., 2025) and DeepSeek-R1 (Guo et al., 2025) as well as proprietary ones, to produce diverse pairs of task-specific specifications and corresponding golden RTL implementations.
-Since the generated RTL may deviate from the original design, we apply an additional filtering stage to remove samples with syntax errors or potential benchmark contamination, identified by high similarity to downstream evaluation benchmarks.
-For modification and debugging tasks, the specification includes a simplified or intentionally faulty version of the golden code, emulating realistic design correction workflows.
-Through this pipeline, we curate a large-scale dataset of 1.7 million specification–RTL pairs that comprehensively cover diverse hardware design scenarios.
-Using this large-scale RTL dataset, we apply supervised fine-tuning (SFT) (Chung et al., 2024) to obtain the RTL-specialized model.
-The resulting model internalizes hardware domain knowledge and RTL semantics, enabling it to generate correct RTL code across diverse design tasks.
-3.2. Reflector: A Frontier Reasoning LLM
-While the Generator produces RTL code grounded in domain knowledge, the Reflector analyzes simulation feedback and guides refinement.
-Frontier generic LLMs excel at interpreting specifications, buggy code, and error reports, making them well suited for this role.
-In ACE-RTL, the Reflector is built upon Claude4-Sonnet to interpret verifiable feedback and generate structured fix strategies.
-At each iteration, the RTL code produced by the Generator is compiled and simulated using iverilog to verify functional correctness.
-When simulation fails, the Reflector is triggered to analyze the feedback.
-Automated scripts invoke the simulator and convert the raw logs into a structured format that captures the error message and the expected versus actual signal behavior.
-Claude4-Sonnet then serves as the Reflector’s reasoning engine: it integrates this parsed information with the design specification and the buggy RTL code to produce a structured diagnostic report.
-The report provides a concise explanation of the root cause together with high-level fix guidance describing how to resolve the primary failure.
-This structured output is then passed to the Coordinator for context evolution.
+These three components are discussed in detail in Sections III-A, III-B, and III-C, respectively.
+In addition, Section III-D introduces a parallel scaling strategy that executes multiple ACE-RTL instances concurrently to speedup convergence efficiency.
+III-A Generator: An RTL-Specialized LLM
+We develop an RTL-specialized model, referred to as ACE-RTL-Generator, that is trained on a large-scale synthetic dataset to capture diverse hardware design patterns and semantics.
+Unlike prior synthetic data generation (SDG) approaches that focus primarily on specification-to-RTL translation, our SDG considers a broader family of tasks that better reflect real-world hardware development workflows, including code editing and debugging.
+These tasks enable the model to learn refinement and error-correction capabilities that are essential for RTL design, particularly when deployed within an agentic workflow.
+Specifically, we construct three categories of training samples under a unified formulation that transforms an input condition into a correct (golden) RTL implementation $C^{\star}$.
+The input condition may consist of a specification alone or a specification paired with an existing design that must be refined or corrected.
+We denote the input as $X$ and define the task as: $X\rightarrow C^{\star}$, where $X$ takes different forms:
+$X=\begin{cases}S&\text{(Spec-to-RTL)}\cr(S,C_{\text{base}})&\text{(Editing)}\cr(S,C_{\text{bug}})&\text{(Debugging)}\end{cases}$ (1)
+Here, $S$ denotes the specification, which includes functional requirements and, depending on the task, may additionally include feature requests or diagnostic information such as signal mismatches.
+$C_{\text{base}}$ represents a partial implementation missing required functionality, while $C_{\text{bug}}$ denotes a faulty implementation containing design errors.
+Both editing and debugging tasks therefore require the model to reason over existing code context and perform targeted transformations, rather than synthesizing designs from scratch.
+In the following, we describe the steps of our SDG pipeline in detail.
+High-Quality RTL Seed Collection
+We begin by collecting approximately five million RTL scripts from public repositories and open hardware projects.
+To obtain a clean pool of seed designs, we apply deduplication, remove machine-generated artifacts (e.g., netlists), filter extreme outliers (30–2,000 lines per module), and validate syntax using Icarus Verilog (iverilog) [31].
+We further perform benchmark decontamination using the Jaccard similarity metric [10], discarding any sample with similarity greater than $0.8$ to evaluation benchmarks.
+After these stages, we retain 157K high-quality RTL scripts, which we use as seed implementations to prompt LLMs for synthesizing diverse RTL problems and corresponding golden solutions.
+Specification-to-RTL Data Construction
+Unlike prior works that rely on chain-of-thought (CoT) reasoning supervision, we intentionally avoid training on CoT traces, since they substantially increase inference latency when the model is deployed within an agentic loop.
+Instead, we use a small set of human-crafted specification-to-RTL examples only as in-context demonstrations to teach frontier LLMs the desired task format and complexity.
+The large-scale synthetic corpus is then constructed from the collected RTL seeds in Step (a): for each seed RTL implementation, we prompt frontier LLMs, including GPT-OSS-120B and DeepSeek-R1, to generate diverse synthetic specification–code pairs conditioned on the seed design.
+Editing and Debugging Data Construction
+Beyond specification-to-RTL generation, we further construct editing and debugging samples to model realistic refinement workflows.
+Starting from the collected RTL seed implementations, which are treated as golden designs for editing and debugging tasks, LLMs are prompted to generate intermediate variants, including simplified versions with missing functionality and buggy versions with injected design errors.
+Simplified variants are created by removing advanced features (e.g., pipeline stages), replacing optimized control logic with baseline implementations, or restricting corner-case handling such as overflow detection or boundary conditions.
+Buggy variants are produced by introducing realistic RTL faults, including incorrect state transitions in finite-state machines, timing or handshake protocol violations, off-by-one counter logic, and related design mistakes.
+Given these intermediate implementations, LLMs then generate corresponding specifications that either describe the required feature extensions (for editing) or provide diagnostic information (for debugging).
+Diagnostic specifications are constructed by analyzing the injected faults in the buggy implementation and generating failure descriptions consistent with the underlying errors.
+The resulting training instances require recovering the correct implementation from incomplete or faulty inputs, encouraging the model to learn iterative refinement and error-correction behaviors.
+Filtering and Quality Control
+All generated samples undergo multi-stage validation to ensure correctness and diversity.
+We first perform syntax validation and benchmark decontamination to ensure structural correctness and prevent data leakage [10] [33].
+We then apply semantic alignment scoring using an LLM-as-Judge framework [11], which evaluates the consistency between specifications and implementations on a 1–5 scale.
+To construct the evaluation framework, we first apply the judge to a subset of training samples and manually verify whether the assigned scores are reasonable.
+If inconsistencies are observed, we refine the evaluation prompt and repeat the process until the scoring behavior is satisfactory.
+The finalized judge is then applied to the full dataset, and samples with scores below 3 are discarded to maintain data quality.
+This strategy provides an efficient way for large-scale SDG while maintaining sufficient quality control.
+Through this pipeline, we curate approximately 1.7 million high-quality training instances spanning diverse RTL coding problems.
+Using this large-scale RTL dataset, we apply supervised fine-tuning (SFT) [6] to obtain the RTL-specialized model.
+The resulting model captures rich hardware domain knowledge and RTL semantics, enabling accurate code generation across diverse design tasks.
+Moreover, training on editing and debugging supervision enables the model to effectively analyze, refine, and correct RTL implementations, making it well suited for deployment within our agentic framework.
+III-B Reflector: A Frontier Reasoning LLM
+While our Generator produces and refines implementations within the agentic loop, effective iteration also requires interpreting simulation outputs and identifying corrective actions.
+We therefore introduce the Reflector, a generic frontier LLM (Claude4-Sonnet) that analyzes simulation feedback and long-context error reports to guide subsequent refinement.
+Processing such complex information benefits from the broader reasoning capability and context capacity of frontier models.
+At each iteration, we invoke iverilog once to compile and execute the candidate RTL against the benchmark test harness.
+When the run fails, the resulting logs are parsed into a structured report capturing the error message together with the expected-versus-actual signal behavior.
+Claude4-Sonnet then serves as the Reflector’s reasoning engine: it combines this structured feedback with the design specification and the current RTL implementation to produce a concise diagnosis and high-level fix guidance.
+This output is then passed to the Coordinator to construct the evolving context for subsequent refinement steps.
 Notably, the entire process depends only on iverilog, a lightweight and open-source simulator, rather than complex proprietary toolchains.
-This design choice makes the system portable, easy to maintain, and readily adaptable to diverse RTL design tasks.
-3.3. Coordinator: A Context Evolution Engine
+III-C Coordinator: A Context Evolution Engine
 The Coordinator in ACE-RTL serves as the memory and control center of the agentic loop, maintaining a dynamic, self-evolving context that guides the Generator in subsequent iterations.
@@
 The Coordinator performs two key functions.
-First, it aggregates debugging history across iterations, tracking which errors were identified, what fixes were suggested, and whether those fixes resolved the issues.
-This structured memory preserves useful progress and prevents regressions, allowing the Generator to focus on remaining failure cases rather than revisiting already-corrected logic.
+First, it aggregates debugging history across iterations, tracking identified errors, suggested fixes, and their outcomes.
+This history is incorporated into the evolving context, preserving useful progress and preventing regressions so that the Generator can focus on unresolved failures.
 Second, the Coordinator monitors progress across iterations.
-If the same error persists over multiple rounds, it recognizes that the current code trajectory may be difficult to repair and triggers a restart mechanism, discarding the current implementation and prompting the Generator to produce a fresh RTL design from the specification.
-This leverages LLM stochasticity to explore alternative implementations that is easier to correct.
-By combining structured history aggregation with an adaptive restart mechanism to construct a self-evolving context, the Coordinator keeps ACE-RTL both stable and flexible.
-It preserves useful debugging insights when progress is made and triggers a controlled reset when stagnation occurs, enabling more efficient progress toward functionally correct RTL code.
-3.4. Parallel Scaling in ACE-RTL
+It uses Claude4-Sonnet to examine the debugging trajectory and determine whether the current search has stalled.
+When the same primary failure persists across several iterations without meaningful improvement, the Coordinator triggers a restart mechanism, discarding the current implementation and prompting the Generator to produce a fresh RTL design.
+This leverages LLM stochasticity to explore alternative implementations that are easier to correct.
+By combining cross-iteration history aggregation with an adaptive restart mechanism, the Coordinator keeps the evolving context informative and flexible.
+It preserves useful debugging insights when progress is made and triggers controlled resets during stagnation, enabling efficient convergence to correct RTL code.
+Notably, from the Generator’s perspective, the self-evolving context resembles an augmented specification that describes failure symptoms or required modifications.
+Because the Generator is trained on both editing and debugging tasks, it is naturally capable of interpreting such context, without requiring additional model adaptation.
+III-D Parallel Scaling in ACE-RTL
 Although the iterative debugging framework in ACE-RTL effectively improves code correctness, it may require a large number of iterations to converge when starting from an unfavorable initial RTL generation.
@@
 If none of the processes achieve functional correctness within the predefined maximum number of iterations, all are stopped automatically.
-This strategy leverages the diversity introduced by LLM stochasticity, as some initial generations naturally require far fewer debugging steps than others.
-By exploring multiple debugging paths in parallel, the system reduces the worst-case iterations needed to reach a correct solution.
-Consequently, the parallel scaling strategy largely reduces the overall iteration count, as confirmed in Figure 6.
-4. Experiment
-We conduct a comprehensive evaluation of ACE-RTL, comparing its performance against $14$ competitive baselines on the CVDP benchmark.
-In addition, we present three case studies illustrating the effectiveness of key components within ACE-RTL.
-Besides, we also demonstrate how the proposed parallel scaling strategy largely reduces the number of iterations required to reach correct solutions.
-4.1. Experimental Setup
-Baselines
-We evaluate ACE-RTL against $13$ SoTA baselines, including latest large open-source models Llama4-Maverick (AI, 2024), DeepSeek-v3.1 (Liu et al., 2024a), DeepSeek-R1 (Guo et al., 2025), Kimi-K2 (Team et al., 2025), and Qwen3-Coder-480B (Yang et al., 2025).
-Frontier proprietary models comprise 4o-mini (Hurst et al., 2024), GPT-5 (OpenAI, 2025), and Claude4-Sonnet (Anthropic, 2025a).
-For RTL models, we include RTLCoder-v1.1, CodeV (Zhao et al., 2025), OriGen (Cui et al., 2024a), CraftRTL, and ScaleRTL.
-Besides, given that ScaleRTL extends ScaleRTL via test-time scaling and achieves performance comparable to prior RTL agentic methods such as VerilogCoder and MAGE, we select it as the representative agentic baseline for evaluation on CVDP.
-Our variants are: (1) ACE-RTL-Generator, our RTL-specialized model, and (2) ACE-RTL, which integrates all components of this work.
-(3) ACE-RTL (Claude4), which uses the same agentic framework as ACE-RTL but replaces our specialized Generator with Claude4-Sonnet.
-CVDP Benchmark
-We adopt the CVDP-v1.0.2 benchmark as our primary evaluation suite, which includes four representative RTL design tasks: code completion (CVDP-cid002), specification-to-RTL generation (CVDP-cid003), code modification (CVDP-cid004), and code debugging (CVDP-cid016).
-For the latter two tasks, the input prompt includes a basic or buggy RTL implementation that is intended to be refined according to the given specification.
-Together, these tasks capture the core challenges faced by RTL designers in practice, covering both generative and corrective design scenarios.
-The benchmark contains $94$, $78$, $55$, and $35$ problems under the four categories respectively, covering a diverse range of hardware domains including arithmetic units, on-chip communication protocols, memory hierarchies, DSP kernels, and control logic.
-We leave testbench generation and subjective evaluation categories in CVDP to our future work.
-Each problem is evaluated across five independent runs.
-We report two complementary metrics: (1) Pass@1, the average success rate across independent runs (Liu et al., 2023b), and (2) Agentic Pass Rate (APR), defined as the ratio of unique solved problems to the total number of problems.
-Since agentic methods inherently perform multiple reasoning and correction iterations, comparing them against standalone LLMs solely using Pass@1 can be misleading.
-Therefore, we introduce APR to better capture a model’s overall problem-solving coverage across multiple trials, providing a fairer assessment of agentic approaches.
-Training and Inference
-ACE-RTL-Generator is finetuned from Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) using 32 compute nodes, each equipped with 8 NVIDIA A100 GPUs.
-We adopt tensor, pipeline, and context parallelism to enable efficient large-scale fine-tuning.
-Training is performed for three epochs with a context window of 32,76832{,}768 tokens and a global batch size of $128$, using a cosine-annealing learning-rate scheduler.
-For inference, ACE-RTL uses five parallel processes under our parallel scaling strategy, with each process limited to 30 iterations and the Generator configured with a temperature of $1.2$.
-The finetuned ACE-RTL-Generator is hosted via vLLM for low-latency API inference (Kwon et al., 2023), while other components leverage Claude4-Sonnet through its official API calls.
-Table 1.
-Functional correctness on the CVDP benchmark.
-Each benchmark reports percentage scores for Pass@1 and Agentic Pass Rate (APR).
-APR is defined as the ratio of unique solved problems to the total number of problems.
-For agentic methods, we omit Pass@1 scores (shown as "—") because their Pass@1 equals their APR.
-We use bold to denote the best results achieved by our methods and underline to mark the strongest baseline results within each CVDP task category.
+This strategy leverages LLM generation diversity to reduce time-to-first-success with a trade-off in total inference-time compute.
+In Section IV-E, we evaluate parallel scaling from both convergence and inference-cost (i.e., token usage) perspectives.
+Our goal is not to claim lower total compute, but to show that parallel scaling provides a favorable speed-cost trade-off.
+IV Experiment
+IV-A Experimental Setup
+Table 1: Functional correctness on CVDP.
+We report Pass@1 and Agentic Pass Rate (APR).
+For agentic methods, Pass@1 is omitted (“—”) since it equals APR.
+Bold indicates our best results, and underline the strongest baseline in each category.
 Type Model CVDP-cid002 Code Completion CVDP-cid003 Spec-to-RTL CVDP-cid004 Code Modification CVDP-cid016 Code Debugging
@@
 DeepSeek-v3.1 $32.34$ $37.23$ $41.79$ $48.72$ $36.73$ $41.82$ $34.86$ $40.00$
-DeepSeek-R1 $34.89$ $\underline{39.36}$ $39.23$ $42.31$ $37.45$ $43.64$ $48.57$ $51.43$
+DeepSeek-R1 $34.89$ $39.36$ $39.23$ $42.31$ $37.45$ $43.64$ $48.57$ $51.43$
 Kimi-K2 $23.40$ $25.53$ $26.67$ $29.49$ $29.09$ $32.73$ $29.71$ $31.43$
@@
 Proprietary models o4-mini $35.10$ $37.23$ $41.56$ $45.45$ $41.48$ $44.44$ $50.00$ $58.82$
-GPT-5 $36.17$ $\underline{39.36}$ $42.31$ $47.44$ $\underline{43.64}$ $45.45$ $\underline{54.28}$ $\underline{60.00}$
-Claude4-Sonnet $\underline{37.94}$ $\underline{39.36}$ $\underline{49.49}$ $\underline{51.28}$ $42.91$ $\underline{49.09}$ $51.43$ $54.29$
+GPT-5 $36.17$ $39.36$ $42.31$ $47.44$ $\underline{43.64}$ $45.45$ $\underline{54.28}$ $60.00$
+Claude4-Sonnet $\underline{37.94}$ $39.36$ $\underline{49.49}$ $51.28$ $42.91$ $49.09$ $51.43$ $54.29$
 RTL-specialized models RTLCoder-v1.1-7B $0.43$ $1.06$ $3.33$ $5.13$ $1.09$ $1.82$ $0.57$ $2.86$
@@
 ScaleRTL-32B $25.32$ $27.66$ $28.97$ $33.33$ $25.82$ $30.91$ $30.86$ $37.14$
-ScaleRTL-32B $29.79$ $35.90$ $32.73$ $40.00$
-Ours ACE-RTL-Generator $\mathbf{39.57}$ $40.43$ $\mathbf{49.74}$ $52.56$ $\mathbf{65.09}$ $67.27$ $\mathbf{56.00}$ $57.14$
-ACE-RTL (Claude4) $\mathbf{80.85}$ $89.74$ $81.82$ $88.57$
+RTL Agents MAGE (Claude4) $44.68$ $53.85$ $50.91$ $60.00$
+MAGE (ACE-RTL-Generator) $\underline{46.81}$ $\underline{55.13}$ $\underline{70.91}$ $\underline{62.86}$
+Ours Generator (standalone) $\mathbf{39.57}$ $40.43$ $\mathbf{49.74}$ $52.56$ $\mathbf{65.09}$ $67.27$ $\mathbf{56.00}$ $57.14$
 ACE-RTL $\mathbf{80.85}$ $\mathbf{96.15}$ $\mathbf{90.91}$ $\mathbf{91.43}$
-4.2. Main Results
-As shown in Table 1, most prior RTL-specialized models struggle on CVDP, despite their strong performance on earlier RTL benchmarks.
-This suggests that models trained mainly on shorter and simpler RTL data generalize poorly to the more complex and realistic design challenges posed by CVDP.
-In contrast, recent frontier proprietary models such as GPT-5 and Claude4-Sonnet show notable improvements and set strong new baselines, benefiting from broader RTL exposure during large-scale pretraining and stronger reasoning capabilities on complex coding tasks.
-Our RTL-specialized model, ACE-RTL-Generator, further advances the RTL coding frontier, outperforming all baselines on most tasks.
+Baselines
+We evaluate ACE-RTL against $14$ SoTA baselines, including frontier open-source models Llama4-Maverick [1], DeepSeek-v3.1 [18], DeepSeek-R1 [12], Kimi-K2 [29], and Qwen3-Coder-480B [32].
+Frontier proprietary models comprise 4o-mini [15], GPT-5 [26], and Claude4-Sonnet [4].
+For RTL models, we include RTLCoder-v1.1, CodeV [36], OriGen [7], CraftRTL, and ScaleRTL.
+In addition, as VerilogCoder relies on a custom Verilog parsing tool that is not readily portable to SystemVerilog designs in CVDP, we adopt MAGE as the RTL agentic baseline.
+To ensure a fair comparison, we use Claude4-Sonnet as the underlying LLM in MAGE, which is the same generic model used in ACE-RTL.
+Our variants are: (1) Generator, our RTL-specialized model; (2) ACE-RTL, the full system integrating the specialized Generator with the Reflector and Coordinator.
+CVDP Benchmark
+We adopt the CVDP-v1.0.2 benchmark as our primary evaluation suite, which includes four representative RTL design tasks: code completion (CVDP-cid002), specification-to-RTL generation (CVDP-cid003), code modification (CVDP-cid004), and code debugging (CVDP-cid016).
+Together, these tasks capture the core challenges faced by RTL designers in practice, covering both generative and corrective design scenarios.
+The benchmark contains $94$, $78$, $55$, and $35$ problems under the four categories respectively, covering a diverse range of hardware domains including arithmetic units, on-chip communication protocols, memory hierarchies, DSP kernels, and control logic.
+We leave testbench generation and subjective evaluation categories in CVDP to our future work.
+Each problem is evaluated across five independent runs.
+We report two complementary metrics: (1) Pass@1, the average success rate across independent runs [20], and (2) Agentic Pass Rate (APR), defined as the ratio of unique solved problems to the total number of problems.
+Since agentic methods inherently perform multiple reasoning and correction iterations, comparing them against standalone LLMs solely using Pass@1 can be misleading.
+Therefore, we introduce APR to better capture a model’s overall problem-solving coverage across multiple trials, providing a fairer assessment of agentic approaches.
+Training and Inference
+ACE-RTL-Generator is finetuned from Qwen2.5-Coder-32B-Instruct [14] using 32 nodes with 8 NVIDIA A100 GPUs each.
+We employ tensor, pipeline, and context parallelism for efficient large-scale training, and adopt sequence packing over the training data to improve token utilization and speed up the training process.
+The model is trained for three epochs with a 32,768-token context window and a global batch size of 128 using a cosine-annealing learning rate.
+The total training process takes approximately 10K GPU-hours, which is a relatively modest cost for training a strong domain-specialized model, especially compared with frontier LLM training, which typically requires multiple orders of magnitude more compute.
+During inference, ACE-RTL runs five parallel processes under our parallel scaling strategy, each limited to 30 iterations with a temperature of 1.2.
+The finetuned Generator is served via vLLM [17], while other components use Claude4-Sonnet through its official API.
+Figure 3: Comparison of frontier generic and RTL-specialized LLMs.
+IV-B Main Results
+Table 1 shows prior RTL LLMs struggle on CVDP despite performing well on earlier RTL benchmarks, suggesting limited generalization from shorter and simpler RTL data to the more complex designs in CVDP.
+In contrast, frontier models such as GPT-5 and Claude4-Sonnet show clear improvements and establish strong new baselines, benefiting from broader RTL exposure during large-scale pretraining and stronger reasoning on complex coding tasks.
+Our RTL-specialized model, ACE-RTL-Generator, further advances the RTL coding frontier, outperforming all standalone model baselines on most tasks.
 It achieves a Pass@1 of $65.09\%$ on Code Modification (cid004), surpassing the SoTA baseline GPT-5 by $21.45\%$.
-This strong performance stems from our high-quality modification data, created by extracting basic RTL code segments and training the model to augment them with missing features, as discussed in Section 3.1.
-By integrating our specialized Generator with the full ACE framework, ACE-RTL achieves the best overall performance on CVDP, with up to a $44.87\%$ APR improvement over SoTA baselines.
-ACE-RTL also outperforms the ACE-RTL (Claude4) variant, showing that an RTL-specialized model is more effective than relying solely on a frontier generic LLM.
-Together, these results validate the effectiveness of ACE-RTL, which iteratively evolves contextual information to guide the specialized model toward accurate RTL coding.
-4.3. Case Studies
-To further demonstrate the effectiveness of our approach, we present three representative case studies.
-The first case showcases how our RTL-specialized model (ACE-RTL-Generator) exhibits a deeper grasp of hardware design principles and canonical RTL patterns compared to generic LLMs.
-The second case illustrates how ACE-RTL-Reflector elevates repeated local failures into functional reasoning, enabling it to diagnose structural errors.
-The third case demonstrates how ACE-RTL-Coordinator identifies unproductive progress and strategically triggers restarts, enabling the system to escape local minima and reach correct RTL implementations.
-Notably, the examples in the latter two case studies require the highest number of iterations to reach correct solutions in our experiments.
-Refer to caption Placeholder.
-Figure 3. Case Study I — RS232 Transmitter Module.
-Case Study I — This case study examines the RS232 Transmitter module, which performs serial data transmission driven by precise baud-rate timing.
-A central component is the baud-rate generator, which produces periodic pulses that step the transmitter state machine through start, data, and stop bits.
-Correct behavior relies on an accurate fractional clock divider to handle the non-integer ratio between the system clock and the target baud rate.
-As shown in Figure 3, Claude4-Sonnet fails to produce a functionally correct implementation due to misunderstanding a fundamental RTL design pattern.
-Its code couples the accumulator and pulse into a single assignment, thereby losing the overflow-based pulse signal and producing unstable timing behavior.
-In contrast, our RTL-specialized model (i.e., ACE-RTL-Generator) correctly implements an accumulator-based divider, updating only the lower bits and using the most significant bit as a stable pulse output.
-This solution reflects an understanding of canonical hardware timing structures, which our model learns from curated RTL training data.
-Placeholder.
-Figure 4. Case Study II.
-Placeholder.
-Figure 5. Case Study III.
-Placeholder.
-Figure 6. Effect of parallel scaling.
-     Case Study II — This case study analyzes the 64b/66b Decoder module, which performs bit-aligned extraction of mixed-mode data streams.
+This strong performance stems from our high-quality modification data, created by extracting basic RTL code segments and training the model to augment them with missing features, as discussed in Section III-A.
+Figure 3 compares standalone models in terms of average CVDP Pass@1 across four categories versus model size.
+In addition to achieving the highest standalone pass rate, ACE-RTL-Generator also lies beyond the Pareto frontier formed by prior baselines, indicating a particularly favorable trade-off between model scale, which is broadly correlated with training cost, and functional correctness.
+For agentic baselines, we evaluate two MAGE variants, using either Claude4 or ACE-RTL-Generator as the underlying RTL generation model.
+In both cases, MAGE achieves only marginal gains over the corresponding standalone model.
+We attribute this mainly to the absence of a self-evolving context that preserves useful learnings from prior iterations.
+Without such accumulated debugging history, the agentic loop has limited ability to build on past failures, making it harder to identify root causes and apply targeted fixes over time.
+A secondary limitation is that MAGE depends on a testbench agent whose outputs can be inaccurate, even when given the golden CVDP test harness, because translating cocotb-based CVDP harnesses into SystemVerilog adds another layer of difficulty on top of the well-known challenge of testbench generation for complex designs [28].
+These results suggest that prior SoTA agentic baselines still struggle to generalize beyond earlier, simpler benchmarks.
+Overall, ACE-RTL achieves the best performance on CVDP, with up to a $41.02\%$ APR improvement (on cid003) over all baselines.
+These results show that combining our RTL-specialized model with iterative context evolution is highly effective for complex RTL tasks.
+IV-C Ablation Study
+Figure 4 reports the averaged APR of several ablated variants of ACE-RTL across the four CVDP categories.
+Relative to the standalone Generator, ACE-RTL w/o Coordinator already yields a clear improvement, confirming the value of simulation-guided iterative refinement.
+The largest gain, however, comes from enabling the Coordinator in ACE-RTL w/o restart, which maintains an evolving context across iterations using debugging history from prior attempts.
+This result shows that preserving cross-iteration history is a key ingredient in our agentic system, allowing the Generator to build on prior progress rather than repeatedly revisiting similar failures.
+Adding the restart mechanism brings another clear improvement over ACE-RTL w/o restart, indicating that preserving the evolving context alone is not always sufficient.
+Some failed trajectories remain difficult to recover through incremental local fixes and instead benefit from regenerating a fresh implementation under the accumulated debugging context.
+This behavior is also illustrated in Figure 7, where restart helps the system escape prolonged stagnation.
+Finally, ACE-Claude4 uses the same ACE framework but replaces our RTL Generator with Claude4.
+Although it achieves strong performance, the full ACE-RTL still performs best overall, demonstrating that agentic context evolution and RTL specialization are complementary.
+Figure 4: Ablation study on ACE-RTL components.
+Figure 5: Case Study I.
+Figure 6: Case Study II.
+Figure 7: Effect of parallel scaling.
+IV-D Case Studies
+To further demonstrate the effectiveness of ACE-RTL, we present two representative case studies.
+The first illustrates how ACE-RTL moves beyond repeated local fixes to uncover the underlying functional mistake in RTL.
+The second illustrates how ACE-RTL identifies unproductive search trajectories and strategically restarts, enabling it to escape local minima and reach a correct RTL implementation.
+     Case Study I — This case study analyzes the 64b/66b Decoder module, which performs bit-aligned extraction of mixed-mode data streams.
 Figure 7 shows that ACE-RTL quickly improves functional correctness but remains stuck for many iterations before passing the final test case.
-ACE-RTL-Reflector identifies this stagnation and analyzes the discrepancies between expected and actual outputs.
-By comparing these discrepancies against the provided specification, the Reflector infers that the decoding logic requires an implicit alignment transformation that is not stated in the specification.
-It then formulates high-level guidance describing this missing behavioral requirement, which directs the Generator to update the extraction logic accordingly, after which the design immediately passes all remaining tests.
-This case illustrates how ACE-RTL-Reflector interprets systematic failure signals to detect specification gaps and elevate repeated local adjustments into understanding the implied design intent, enabling the system to resolve design-level issues that cannot be addressed through isolated fixes.
-Case Study III — This case study analyzes the Clock Jitter Detection module, which evaluates clock stability by comparing measured edge intervals against a timing threshold.
+Our Reflector identifies this stagnation and analyzes the discrepancies between expected and actual outputs.
+By comparing these discrepancies against the provided specification, the Reflector infers that the decoding logic requires an implicit alignment transformation that is not clearly stated in the specification.
+It then formulates high-level guidance describing this behavioral requirement, which directs the Generator to update the extraction logic accordingly, after which the design immediately passes all remaining tests.
+Case Study II — This case study analyzes the Clock Jitter Detection module, which evaluates clock stability by comparing measured edge intervals against a timing threshold.
 As shown in Figure 7, ACE-RTL makes early progress but soon enters a long plateau where all attempts fail.
@@
 Although the Generator produces many syntactic variants, they all retain this early-trigger issue and therefore cannot satisfy the required timing semantics.
-ACE-RTL-Coordinator recognizes this stalled trajectory by observing that many consecutive iterations exhibit the same assertion pattern with no meaningful improvement.
+Our Coordinator recognizes this stalled trajectory by observing that many consecutive iterations exhibit the same assertion pattern with no meaningful improvement.
 Consequently, the Coordinator triggers a restart and distills high-level insights from the failed attempts, such as the need to validate the first measurement interval before performing detection.
 It then instructs the Generator to regenerate the design from the specification using these clarified constraints.
-The restart produces a simpler and more interpretable implementation that shifts the logic toward proper interval validation.
-After two restarts, the Generator introduces a validity flag that delays jitter detection until the first full measurement interval has been observed, which immediately restores correctness.
-This case shows how ACE-RTL-Coordinator identifies prolonged stagnation, extracts useful design insights, and uses restart mechanism to move the system toward correct RTL implementations.
-4.4. Efficacy of Parallel Scaling
-Figure 6 illustrates the impact of the proposed parallel scaling strategy on reducing the iteration count required for convergence across multiple CVDP categories.
-The results clearly show that launching multiple independent ACE-RTL instances in parallel significantly accelerates the debugging process without compromising correctness.
-Specifically, the average iteration count drops from $11.33$ to $3.95$ for cid002, from $11.25$ to $4.23$ for cid003, from $9.25$ to $3.75$ for cid004, and from $13.36$ to $4.37$ for cid016, yielding speedups of $2.87\times$, $2.66\times$, $2.47\times$, and $3.06\times$, respectively.
-This improvement demonstrates the efficacy of the parallel scaling design: by exploring diverse debugging trajectories derived from different initial generations, ACE-RTL avoids being constrained by any single unfavorable starting point.
-Instead, at least one parallel process typically converges rapidly, allowing other trajectories to be terminated early.
-Consequently, the system achieves nearly a $3\times$ runtime reduction to explore a correct solution.
-4.5. Results on VerilogEval
-Table 2. Pass rate results on VerilogEval-Human-v2 — ACE-RTL-G denotes the standalone ACE-RTL-Generator model.
+The restart yields a simpler implementation that properly validates the measurement interval.
+After two restarts, the Generator introduces a validity flag that delays jitter detection until the first full interval is observed, restoring correctness.
+This example illustrates how the Coordinator detects stagnation and uses a restart mechanism to guide the system toward a correct RTL implementation.
+Table 2: Impact of PS (parallel scaling) on convergence speed and inference cost, averaged over the four CVDP categories.
+Setting APR $\uparrow$ Iter. $\downarrow$ Speedup $\uparrow$ Rel. Token $\downarrow$
+w/o PS 89.84 11.30 $\times 1.00$$\times$
+[…]
+IV-E Efficacy of Parallel Scaling
+[…]
+IV-F Results on VerilogEval
+Table 3: Pass rate results on VerilogEval-Human-v2 — ACE-RTL-G denotes the standalone ACE-RTL-Generator model.
 Pass@1 APR
-CraftRTL Claude4-Sonnet ACE-RTL-G ScaleRTL VerilogCoder ACE-RTL
-$68.0$ $73.0$ $\mathbf{73.8}$ $93.6$ $94.2$ $\mathbf{95.5}$
+CraftRTL Claude4 ACE-RTL-G VerilogCoder MAGE ACE-RTL
+$68.0$ $73.0$ $\mathbf{73.8}$ $94.2$ $\mathbf{95.5}$ $\mathbf{95.5}$
 For completeness, we also evaluate our approach on VerilogEval-Human-v2.
-To ensure fair and rigorous comparison, we use Pass@1 for standalone models and APR for agentic methods.
-As shown in Table 2, ACE-RTL-G outperforms both the strong RTL model CraftRTL and the frontier model Claude4-Sonnet, demonstrating the strength of our RTL-specialized generator.
-Moreover, ACE-RTL achieves the highest pass rate among all agentic baselines, surpassing ScaleRTL and VerilogCoder.
+We use Pass@1 for standalone models and APR for agentic methods.
+As shown in Table 3, ACE-RTL-G outperforms prior RTL LLMs such as OriGen and CraftRTL.
+Moreover, ACE-RTL achieves the highest pass rate among all agentic baselines.
 These results show that ACE-RTL is not overfitted to CVDP but instead captures transferable RTL reasoning and coding capabilities, maintaining strong performance on earlier RTL benchmarks.
-5. Conclusion
-This work introduces ACE-RTL, a framework that integrates an RTL-specialized model with ACE for accurate RTL code generation.
-ACE-RTL achieves SoTA results on multiple RTL benchmarks, surpassing both frontier proprietary and RTL-specialized models.
+V Failure Analysis and Future Work
+Although ACE-RTL achieves strong results on CVDP, it does not yet reach a $100\%$ pass rate.
+We observe two main sources of remaining failures.
+First, some datapoints require fine-grained temporal diagnosis, where textual simulation feedback alone is insufficient to localize the true timing-related bug.
+Second, some failures arise from ambiguity in the specification, or from subtle mismatches between the intended design behavior and the behavior emphasized by the test cases.
+These cases require not only debugging against observed failures, but also reasoning about the true design intent described in the specification.
+One promising direction for addressing these failures is to incorporate waveform information, which can reveal temporal mismatches more directly than textual logs.
+However, we intentionally exclude waveform information in this work to evaluate whether ACE-RTL can solve RTL tasks primarily through RTL knowledge acquired via domain-specific training and specification-grounded reasoning, using only limited non-waveform simulation feedback.
+Although waveform-driven debugging often boosts the pass rate of LLM agents, it also risks overfitting to observed test behavior while underutilizing the specification itself, especially in cases with specification ambiguity or specification-test mismatch.
+We therefore view waveform information not as a substitute for specification-grounded reasoning, but as a potentially useful auxiliary signal that should be integrated into ACE-RTL in a careful and principled way to avoid overfitting to test cases.
+We leave such integration as an important future direction.
+VI Conclusion
+This work introduces ACE-RTL, a framework that integrates an RTL model with ACE and achieves SoTA results on multiple RTL benchmarks.
+An interesting future direction is to automate testbench generation and explore waveform-enhanced agentic reasoning.